# Vulnerabilities that arise from poor governance in Distributed Ledger Technologies

### v1

@@ 5. Missing properties
This section outlines the number of currently disregarded good governance properties. Many DLT governance protocols are currently failing to meet standards of good governance in their implementations. Some of these include: 1. Binding Elections: Many governance protocols host their voting on off-chain platforms, meaning the results are not binding. Similarly, some schemes allow participants to vote that are not responsible for implementing the results.
@@
6. Verifiability: Voters should be able to verify the election outcome reflects their votes.
@@
The absence of these governance properties results in a number of possible exploits. We further outline these in the subsequent section. In current DLTs, achieving all the above properties is not possible, since some are mutually exclusive. We encourage practitioners to consider the context specific security and performance needs of their DLT and select a compromise appropriately. We proceed by outlining which properties are mutually exclusive and why that is the case.
@@
Non-purchasable voting power and permissionless access to the DLT: This is currently the most prevalent issue in DLT governance. DLTs were designed to avoid having to trust central authorities in order to deliver a service. In order to achieve this whilst being protected against resource attacks, they must still offer some protection against Sybil attacks. In other words: because DLTs aim to be decentralised, they must avoid scenarios where any actor can access the network and attack it. Therefore, access to the DLT must be regulated. Currently, to avoid doing this by having to again rely on a centralised trust authority to grant access credentials, access is regulated through a proxy for wealth. The wealthier amass more voting power as a consequence. De-facto, this re-introduces centralisation in the DLT. We advocate for access control mechanisms that do not prioritise the wealthier, but are based on other criteria. An example is the Tree-Proof-of-Position mechanism presented in (Kharman et al., 2023). This work presents a decentralised class of algorithms that allow agents to prove they are in a given position in a collaborative manner. A scheme like this may be used to regulate access to DLTs, such that only agents with a valid proof of position may be granted participation rights.
@@
6. Vulnerabilities
6.1. Exploits leveraging public votes
The right to a private vote is recognized as a human right as per Article 21.3 of the Universal Declaration of Human Rights (UN General Assembly, 1948) in 1948. In the absence of it, voters may be coerced, bribed or sell their votes. Secret ballots are congruent to freedom, they ensure voters can express their will free of external influences and pressures. This is no different in DLT governance. The absence of secret ballots is responsible for a suite of vulnerabilities that we now proceed to outline.
6.1.1. 1. Voter Coercion
When ballots are public, everyone can know how a voter voted. This provides verifiability: everyone can check that the election outcome is correct and that only those allowed to vote voted. However, this forgoes freedom. Free will cannot be protected when a voter faces the reality of not abiding with their coercer’s threats (Lever, 2007). The coercer can easily check if their victim complied. Voters may chose to sell their votes even in the absence of coercion. In some cases, DLTs provide a unique set of conditions where the most benefitial strategy for some voters is to sell their vote. Other works have identified this issue in DAO governance (Daian et al., 2018), (Austgen et al., 2023). We highlight that this problem is prevalent in DLT governance too.
We stress, at current time of writing, none of these solutions offer secret ballots. No existing solution can be shown to satisfy formal notions of ballot secrecy, because the ledger is public. Some have attempted to provide increased levels of privacy, but this privacy is short lived, and often compromised to achieve verifiability. An example is Snapshot’s recently added feature of ‘shielded voting’, wherein votes remain private only until the end of the election. Then, ‘when voting ends, all votes are revealed - as well as who voted for what’  [^3] Snapshot adds ‘Shielded Voting’. This still enables voter coercion and vote selling.
6.1.2. 2. Disincentive to vote against whales
Both users and miners vote with their feet on a DLT. When a proposal arises for a protocol update, both must agree for it to be implemented. Users may favour a certain proposal, but miners can refuse to service transactions following it. Similarly, if miners favour a proposal that users disagree with, the users may simply stop sending transactions on that chain, and miners will earn no fees. Let us consider a voter with much larger voting power than the rest a whale. An example could be a user with a large share of tokens or a collective with high mining power (a large mining pool) in the DLT. When a whale votes for a certain proposal, other agents are incentivised to follow suit. Otherwise a fork may arise, and eventually one of the two will outgrow the other and become the main chain. Miners processing transactions on the losing fork will not receive their reward. Consequently users’ transactions will not get approved, and thus they too will eventually leave the fork. Because whales hold greater voting power, the fork they vote for will be the winning one with greater likelihood. Less influential voters will minimise their chance of losses by following the behaviour of whales, which is public, on-chain. This creates a herding effect, and a disincentive to vote true to one’s beliefs, if these oppose those of a whale’s.
6.1.3. 3. Voter Information Asymmetry
‘Pump and dumps’ become easy: a proposal may be submitted that stands to inflate token prices. A whale holding many of these tokens may informally communicate that they will vote for this proposal (i.e.: via Twitter, Discord or other similar channels). Less influential voters begin to cast their votes in agreement (we explain above how there is no incentive to vote against the views of a whale). Token prices begin to surge, as it is expected for this proposal to pass. Participants rush to buy more of these tokens. All the meanwhile the whale has not cast their vote yet. Just before the end of the election, the whale sells their tokens, casts their vote, and thus crashes the token price. They have made a profit by cashing out, exploiting the voter information asymmetry that public votes provide, tanking the token price in the process. The public votes allows all participants in the DLT to behave in agreement, creating a herding effect. It is this behaviour that the whale can then exploit for their own profit. Private votes prevent this scenario.
Another simpler exploit consists of an attacker purchasing enough voting power to approve a malicious proposal in their favour. Public voting facilitates this attack: the attacker knows exactly how much voting power they need to acquire to overturn an election in their favour. These attacks have been witnessed in DAOs a number of times  [^4] Build DAO’s hostile governance takeover attack, Feb 2022,  [^5] Beanstalk cryptocurrency project robbed after hacker votes to send themselves \$182 million,  [^6] Sanctioned Tornado Cash DAO governance heisted by hacker, and are plausible in DLTs too. As shown in (Wang et al., 2020) and (Gencer et al., 2018), miners’ voting power (hash rate distribution) is highly centralized. In some cases, the voting power held by one agent can suffice to fork a network. In others, malicious users may leverage flash loans [^7] Loans written in smart contracts that enable participants to quickly borrow funds without the need for collateral (Chainalysis, 2023). to rapidly obtain sufficient tokens, submit a proposal to their benefit, cash out and repay the loan, whilst still making a profit. Miners are incentivised to service this proposal: they stand to gain high transaction fees. Indeed, the behaviour of whales is known to affect the price of a token (Manahov, 2022): they have the power to send large scale transactions that may incentivize miners to act against the consensus protocol (Liao and Katz, 2017), (Daian et al., 2020).
6.2. Exploit leveraging non-binding elections
We note that, although informal voting channels may exist, these are in fact nothing other than a polling mechanism, because the only binding vote on a DLT protocol update is in fact the transactions sent by users and approved by miners on-chain, following either one protocol or another. All other off-chain so-called voting platforms are operating under the assumption that miners and users will behave on-chain in the same way that they expressed they would off-chain. The New York Agreement is a clear example of how off-chain voting is not binding and how this assumption may not always hold [^8] Bitcoin’s New York Agreement (Hunter, 2022). Anyone can vote, but if the agents responsible for implementing the election outcome disagree with the outcome, they have no obligation to implement it, even if the election was fair, free and correct. This is a vulnerability, as stakeholders may subvert an election outcome by refusing to comply. To prevent this, DLT governance protocols must be binding. To ensure this, elections must be carried out on-chain. On-chain voting brings a suite of challenges with it, which we outline in section 6.3.2.
6.3. Exploits leveraging purchasable voting power
6.3.1. 1. Plutocracy
1-dollar-1-vote mechanisms and any variant of these where voting power may be purchased will converge towards granting power to the most wealthy. Any governance mechanism that allows users to purchase voting power will inherently converge to plutocracy and centralise over time (Reijers et al., 2021). Whilst this issue is well known, DLTs continue to implement token based voting, for lack of a better alternative for a decentralised Sybil-protection system. DLT practitioners must compromise and choose which of the alternatives suits their application needs best: permissionless voting does not rely on a central trusted service to prevent Sybil attacks, and is often implemented by providing voting rights proportional to wealth or computational power. However, it is known to centralise voting power in the hands of the wealthy. Permissioned voting introduces reliance on a centralised trusted authority to issue voting rights, or requires a practical and functional decentralised identity management solution. This permissioned alternative is often implemented through reputation based voting schemes, soulbound tokens, NFTs, or a Know Your Customer solution, checking the validity of participants’ identity.
6.3.2. 2. On-Chain Voting
6.4. Other vulnerabilities
6.4.1. 1. Bribe Automation
This exploit is a consequence of the following properties being absent: ballot secrecy, non-purchaseable voting power and fee-less voting. In DLTs where mining is competitive, if one mining pool gains a transaction, the other do not, so we can model the mining pools to be explicitly distinct from each other, and the smallest mining pool is of size 1 agent. Public votes make it evident to see how a mining pool voted by simply looking up the on-chain transactions. Less influential miners should either align their vote to those of the largest mining pools, or risk having their transactions end up in a fork and not receive the reward for the work they performed in including it. It is therefore easy for a briber to identify miners of less influence and bribe them to ‘vote’ for their desired protocol. The small sized mining pools will make a profit, and the briber can identify how many miners they must bribe to succeed in implementing their desired protocol. As the authors in (Austgen et al., 2023) showed, paradoxically, the more decentralised the network, the easier it is to create a systematic bribing cartel. Small sized miners are cheaper to bribe than large mining pools. Indeed, they even provide a protocol to automate this vote buying in DAOs. This can be re-purposed to buy a miner’s vote in a DLT. The bribe can be paid out as soon as the smart contract detects a transaction following the desired protocol of the briber.
6.4.2. 2. Submitting Malicious Protocol Updates:
Malicious protocol updates are not infrequent, and not only in DAOs or DLTs  [^9] Backdoor found in Library update, used by major Linux Distributions. Some go undetected due to low participation rates in governance, in others, the protocol update is forcibly approved by a whale, and in other occasions, the malicious protocol is hidden or not evidently detectable by the community. The consequences of this can be dire: double spending attacks, delaying governance outcomes, creating new tokens or loss of funds are amongst the many attack vectors possible.
It is not feasible to characterise all the ways in which malicious protocol updates may be submitted. This issue is application specific. Governance is only one of the components necessary in the DLT to prevent these exploits, but a robust governance protocol does not guarantee protection against malicious protocol updates. Access control mechanisms, security and trust assumptions are other components that must be taken into consideration. We encourage DLT design practitioners to make choices in their governance protocol that will alleviate the risk of this exploit. Namely, removing incentives to submit malicious proposals, or to comply with bribes to implement them.
@@ 7. Conclusion
We call for the DLT research and practitioners community to consider the matter of governance carefully. Current governance models make DLTs privy to a number of undesirable scenarios. Our strong encouragement is for practitioners to adopt well established notions of security in their own governance protocols, borrowing from the research in social choice theory, cryptography and e-voting schemes. We re-iterate notions of good governance properties, and present a collection of vulnerabilities that arise in the absence of them. We point practitioners towards exiting solutions that may serve as a starting point to begin to deliver good governance properties.
@@ Acknowledgements.
Aida Manzano Kharman thanks IOTA Foundation for funding of her PhD studies. The authors thank Dr. Pietro Ferraro for his valuable advice.

### v2

@@ 8.5. Missing properties
1. Binding Elections: some governance protocols carry out their decision-making process on off-chain platforms (Foundation, 2015; Community, 2011), meaning the results are not binding. Similarly, some schemes allow participants to vote that are not responsible for implementing the results (Community, 2011).
@@
In current DLTs, achieving all the above properties is not possible, since some are mutually exclusive. We encourage practitioners to consider the context specific security and performance needs of their DLT and select a compromise appropriately. We proceed by outlining which properties are mutually exclusive and why that is the case.
@@
Non-purchasable voting power and permissionless access to the DLT: this is currently the most prevalent issue in DLT governance. DLTs were designed to avoid having to trust central authorities in order to deliver a service. In order to achieve this whilst being protected against resource attacks, they must still offer some protection against Sybil attacks [^19] A Sybil attack is one where an agent easily obtains multiple identities in the network to gain unfair advantages (Mohaisen and Kim, 2013).. In other words: because DLTs aim to be decentralised, they must avoid scenarios where any actor can access the network and attack it. Therefore, access to the DLT must be regulated. Currently, to avoid doing this by having to again rely on a centralised trust authority to grant access credentials, access is regulated through a proxy for wealth. The wealthier amass more voting power as a consequence. De-facto, this re-introduces centralisation in the DLT. We advocate for access control mechanisms that do not prioritise the wealthier, but are based on other criteria. An example is the Tree-Proof-of-Position mechanism presented in (Kharman et al., 2023). This work presents a decentralised class of algorithms that allow agents to prove they are in a given position in a collaborative manner. A scheme like this may be used to regulate access to DLTs, such that only agents with a valid proof of position may be granted participation rights.
@@ 9. Conclusion
This paper underscores the critical role of governance in ensuring the security, fairness, and sustainability of Distributed Ledger Technologies (DLTs). By identifying and categorising vulnerabilities that arise from poor governance, we have shown how these systems, despite their promise of decentralisation and transparency, remain susceptible to exploitation. The outlined threats—ranging from coercion and vote-buying to centralization of power—highlight the urgent need for robust governance protocols that align with established principles of fairness, security, and accountability.
Our key contributions include:
  • A taxonomy of governance properties that defines the minimum requirements for ensuring good governance in DLTs.
  • A comprehensive analysis of vulnerabilities stemming from governance deficiencies and their implications for stakeholders.
  • Guidance on technical solutions, bridging insights from cryptography, social choice theory, and e-voting systems to propose actionable pathways for practitioners to implement good governance properties.
This work is necessary because the stakes are high: as DLTs increasingly intersect with public policy, financial systems, and societal infrastructures, the absence of proper governance safeguards risks undermining trust, enabling scams, heists, hacks, centralisation of power and rug pulls. By addressing these vulnerabilities, we not only protect the integrity of DLTs but also uphold their potential to serve as tools for social good.
Future work should focus on:
  1. (1)
    Developing scalable technical solutions for implementing governance properties. Particular focus should be placed on alleviating the severity of properties that are mutually exclusive for good governance. A key point of improvement should be developing permissionless access to DLTs that does not enable purchasable voting power, and delivering fee-less participation.
  2. (2)
    Empirical evaluation frameworks of governance models should be further developed to measure their efficacy in real-world scenarios.
  3. (3)
    Interdisciplinary collaboration to integrate insights from regulatory, ethical, and socio-technical perspectives into governance design. It is crucial to include clear and tangible definitions of good governance in new policies and regulation. Otherwise, it will be impossible to hold projects liable and prevent undesirable vulnerabilities.
  4. (4)
    Standardization efforts to define and promote universally accepted frameworks for good governance in DLTs.
We call upon the DLT community—researchers, developers, users, and policymakers—to prioritize governance as a fundamental component of system design. Only through collective effort and innovation can we ensure that DLTs fulfill their potential as equitable and secure platforms for a decentralized future.
@@ Acknowledgements.
Aida Manzano Kharman gratefully acknowledges the IOTA Foundation for funding her PhD studies. William Sanders was employed by the IOTA Foundation during the period in which the research for this paper was conducted. The authors would like to thank Professor Robert Shorten and Dr. Pietro Ferraro for their insightful advice and valuable contributions.